# Aligning Large Language Models by On-Policy Self-Judgment

### v2

@@ 1 Introduction
In contrast, a line of research Rafailov et al. (2023); Azar et al. (2023) proposes discarding on-policy learning and optimizing from preference orders in static datasets without RMs (i.e., offline learning), whereas other studies propose constructing datasets with responses sampled from the policy and labeled by separated evaluators (i.e., off-policy learning) Zhao et al. (2023); Liu et al. (2023); Xu et al. (2023); Gulcehre et al. (2023); Dong et al. (2023). However, offline learning can lead to sub-optimal results due to the lack of exploration Mediratta et al. (2023), and off-policy learning could cause degeneration if an appropriate replay buffer strategy is not used Zhang and Sutton (2017).
@@ 3.1 Judge Model
Using LLMs to evaluate responses of another LLM, LLM-as-a-judge, is shown promising Zheng et al. (2023); Bai et al. (2022b); Kim et al. (2023). Inspired by these studies, we leverage the generative pairwise evaluator, which we refer to as Judge Model (JM), for aligning LLMs with human preference Ethayarajh et al. (2023); Zhao et al. (2023); Liu et al. (2023). Unlike an RM, which produces a scalar score for a single response, JM simply chooses the better response between the two responses. Specifically, the JM $\pi$ is trained by maximizing the log-likelihood of the choice token $\mathcal{J}\in\{\mathcal{A},\mathcal{B}\}$ corresponding ground-truth preference label for a given judgment template $\mathcal{C}$. Figure 3 illustrates an example of a judgment template.
Figure 3: An example of a judgment template $\mathcal{C}$. The judgment template asks which of the two responses is better for a given prompt and requests to select the choice token $\mathcal{J}\in\{\mathcal{A},\mathcal{B}\}$ corresponding to the better response. Optionally, a principle for the judgment can be added to the judgment template, and the rationale can be included in the target sequence for training.
@@
One advantage of JM compared to the conventional method with RM is that the JM can adapt the judgment template $\mathcal{C}$ according to multiple aspects of human preferences, reflecting diverse principles Sun et al. (2023); Cui et al. (2023). Also, the generative nature of JM makes it easy to utilize rationale, the justification for the judgment in natural language, for more precise judgments, similar to the observations on instruction-tuning of LLMs Mukherjee et al. (2023). For a given principle $p\in\mathcal{P}$ where the $\mathcal{P}$ is a principle set, we devise different judgment templates $\mathcal{C}_{p}$ for each $p$. Also, we can optionally include a rationale $\mathcal{R}$ in the target sequence for the judgment task when it is available. The JM first produces a choice token $\mathcal{J}\in\{\mathcal{A},\mathcal{B}\}$ which corresponds to the ground-truth label of the better response, then rationale $\mathcal{R}$ is followed. Formally, $\pi(\mathcal{J},\mathcal{R}\mid\mathcal{C}_{p}(x,y_{w},y_{l}))$. As a result, principle-aware judgment can be conducted by simply adjusting the judgment template.
@@
Since JM is trained by JSFT, it can judge its own response by playing the roles of the policy and judge: this is the intuition behind the name Self-Judge. As a policy, the JM first samples responses $y_{a},y_{b}$ for a given $x$. Then, the JM, as a judge, chooses a better response between the two responses, $\pi(\mathcal{J}\mid\mathcal{C}(x,y_{a},y_{b}))$. More precisely, it averages the likelihoods of corresponding choice tokens, $\mathcal{J}$, utilizing a position-swapped judgment template to mitigate position bias Chiang et al. (2023). From the relative likelihoods of these choice tokens for each response, a pseudo-preference triplet label $(x,\hat{y}_{w},\hat{y}_{l})$ for the self-training and self-rejection can be deduced. For the principle-aware judgments, we can check the winning rate across principles or the mean likelihood of choice tokens across principles when a tie happens.
@@ 3.2 Self-Training by On-Policy Judgment
Ethayarajh et al. (2023) uses JM’s predicted likelihood of label tokens as the reward for on-policy evaluation in RLHF framework, comparing with a blank baseline response 222huggingface.co/stanfordnlp/SteamSHP-flan-t5-xl. This approach has a significant limitation; the likelihood inferred by a language model cannot be regarded as confidence without calibration Zhou et al. (2023); Zhu et al. (2023). Conversely, to leverage JMs properly, we adopt objectives that perform optimizations on preference orders since these objectives do not require pointwise scalar scores Rafailov et al. (2023); Zhao et al. (2023); Azar et al. (2023). In our framework, we regard the reference policy $\pi_{\text{ref}}$ as a judge to evaluate the preference order between generated samples from current policy $\pi_{\theta}$, as follows:
@@ 3.3 Self-Rejection by Tournament
However, judging all the possible comparison pairs to get average winning rates requires $\mathcal{O}(N^{2})$ forward passes for (${N\choose 2}$ comparisons. Thus, we adopt the tournament Zhao et al. (2023); Liu et al. (2023) for rejection sampling at inference time. We construct a tournament tree whose leaf nodes are sampled responses, and the non-leaf nodes are chosen by the winner on judgment between the child nodes. Since the tournament tree has less than $N-1$ non-leaf nodes, we can find the best response by $\mathcal{O}(N)$ forward passes, the same as the Best-of-$N$ sampling with a separated RM.
@@ 4.2 Baselines
We choose DPO Rafailov et al. (2023) as an offline learning baseline and use DPO objective for self-training in Self-Judge. We include RSO Liu et al. (2023) with the DPO objective as an off-policy learning baseline, which utilizes the JM as a separate evaluator. We include baselines which utilize RMs, ReST Gulcehre et al. (2023) and RAFT Dong et al. (2023) for off-policy approach and RLHF Ziegler et al. (2020) for on-policy approach. We choose LLaMa-2-7B Touvron et al. (2023) as a base model for all experiments.
@@ 5.2 Analysis of Judge Models
Table 2: Prediction accuracy on test split of HH-Helpful and winning rates on AlpacaEval using JM as a judge or a policy. JSFT improves not only the performance as a policy but also the performance as a judge of JM.
@@
In Footnote 4, we observe that RLHF with JM as an evaluator results in a lower performance compared to conventional RM producing scalar rewards, which implies that the choice token likelihood obtained from a judgment could not be regarded as a pointwise preference score on the Bradley-Terry model Bradley and Terry (1952). Table 4 shows that on-policy learning outperforms offline and off-policy approaches using a single JM on self-training. Therefore, we can conclude that on-policy learning with preference orders, such as DPO objective, is the best strategy for JM in order to leverage JM’s superior performance on judgments compared to the RM.
@@
Table 3: Prediction accuracy with different evaluator types. With JMs, we regard the likelihood of choice token comparing with chosen response $y_{w}$ as a reward [^4] We found that using the blank response fails on the assessment of response since the reward is always close to 1.. JM’s token likelihoods are not appropriate for the pointwise reward function in RLHF.
@@ 5.3 Effects of Principle and Rationale
Figure 4: Winning rates, average response lengths, and 4-gram repetitions on AlpacaEval according to the number of sampling ($N$) for self-rejection on JM and JM-PR after self-training. Even though LLM-as-a-judge tends to favor verbose responses Zheng et al. (2023), JM-PR reliably improves winning rates as $N$ increases, with smoother increments of response lengths and lower repetitions compared to JM.
@@ 6 Related Work
From preference orders in the static dataset, DPO Rafailov et al. (2023) optimizes LLMs by implicit rewards without a separated RM. IPO Azar et al. (2023) proposes a modified objective using an unbounded preference mapping function to mitigate overfitting on deterministic preferences in the dataset. PCO Xu et al. (2023) utilizes cringe loss for optimization, which considers the token-level likelihood of rejected samples as contrastive training. SPIN Chen et al. (2024) performs iterative training considering the distribution gap between SFT datasets and the model’s responses as preference orders. Self-Rewarding Language Models Yuan et al. (2024) train LLMs to generate scores for a given response by chain-of-thought reasoning to construct preference datasets by self-generated responses. All these approaches differ from our work in that they do not perform on-policy learning.
@@
The generative pairwise evaluator, which we refer to as JM, has been utilized in previous approaches to alignment learning. ILF Scheurer et al. (2023) selects the response that reflects human-requested feedback through JM. SLiC-HF Zhao et al. (2023) constructs a static preference dataset with responses obtained from the SFT model ordered by JM. RSO Liu et al. (2023) approximates the optimal policy of the RLHF objective with rejection sampling through JM’s choice token likelihood. OAIF Guo et al. (2024) adopts pre-aligned LLMs to perform the pairwise judgment task for on-policy learning, not by fine-tuning to JM. All of these approaches focus on utilizing a separate evaluator and do not address self-improvement.
@@ 7 Discussion
Recent works such as Self-Rewarding Language Models Yuan et al. (2024) and OAIF Guo et al. (2024), which are optimized with pairwise preferences, have intersections with Self-Judge in aspects of self-training and on-policy sampling. However, Self-Rewarding Language Models could not provide on-policy feedback because evaluations are performed through chain-of-thought reasoning. Also, OAIF does not address self-training since it employs a pre-aligned language model. Table 7 illustrates the difference between these approaches compared to Self-Judge.
@@
Table 7: Comparison between Self-Judge and recent concurrent works on preference alignment learning.
@@ Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) Grant funded by the Korea government(MSIT), Artificial Intelligence Graduate School Program, Yonsei University, under Grant 2020-0-01361. The authors thank the members of NAVER Cloud and Yonsei University for their constructive comments. We are also grateful to Seungju Han for valuable discussions and helpful feedback on earlier drafts of this paper.
@@ A.5 Prompt Templates
[…]
@@
HH-Helpful Below are a series of dialogues between a human user and an AI assistant. The AI tries to be helpful, harmless, and honest. UltraFeedback Below are a series of dialogues between a human user and an AI assistant. The AI tries to be helpful, honest, truthful and follows the instructions from human.
@@
[…]
@@
<|system|> {default system message} <|user|> Which is the better response to be an assistant who is helpful, honest, truthful and following the given instruction from user? Instruction: ------- {prompt} ------- Response A: ------- {response A} ------- Response B: ------- {response B} ------- Please choose either A or B. <|assistant|> Sure! The option which is more helpful, honest, truthful and following the given instruction from user would be Response {A or B}
@@
Helpfulness Under the principle of ’helpfulness’, the assistant should provide users with accurate, relevant, and up-to-date information, ensuring that the content is positive, interesting, engaging, educational, and helpful. Honesty Under the principle of ’honesty’, the assistant should be honest about whether it knows the answer and express its uncertainty explicitly. The assistant should confident on questions it knows well and be modest on those it is unfamiliar with using weakeners such as ’I guess’, ’I suppose’, ’probably’, and ’perhaps’ to express uncertainty. Instruction Following Under the principle of ’instruction following’, the assistant should align the output with intent of instruction, by understanding the task goal (intended outcome) and restrictions (text styles, format or designated methods, etc.). Truthfulness Under the principle of ’truthfulness’, the assistant should answer truthfully and be faithful to factual knowledge as well as given contexts, never making up any new facts that aren’t true or cannot be grounded in the instruction.
@@
<|system|> {principle-aware system message} <|user|> Which is the better response for an assistant when following the principle of ’{principle}’ for a given instruction? Instruction: ------- {prompt} ------- Response A: ------- {response A} ------- Response B: ------- {response B} ------- Please choose either A or B according to the principle of ’{principle}’. <|assistant|> Sure! The option which is better guided by the principle of ’{principle}’ would be Response {A or B}
@@
<|system|> {principle-aware system message} <|user|> Which is the better response for an assistant when following the principle of ’{principle}’ for a given instruction? Instruction: ------- {prompt} ------- Response A: ------- {response A} ------- Response B: ------- {response B} ------- Please choose either A or B and evaluate each response according to the principle of ’{principle}’. <|assistant|> Sure! Here is the decision of better response and evaluations of each response, guided by the principle of ’{principle}’. Decision: Response {A or B} Evaluation of Response A: {rationale A} Evaluation of Response B: {rationale B}

### v3

@@ 1 Introduction
In contrast, a line of research Rafailov et al. (2023); Azar et al. (2023) proposes discarding on-policy learning and optimizing from preference orders in static datasets without RMs (i.e., offline learning), whereas other studies propose constructing datasets with responses sampled from the initial policy and labeled by separated evaluators (i.e., off-policy learning) Zhao et al. (2023); Liu et al. (2023); Xu et al. (2023); Gulcehre et al. (2023); Dong et al. (2023). However, offline learning can lead to sub-optimal results due to the lack of exploration Mediratta et al. (2023), and off-policy learning could cause degeneration if an appropriate replay buffer strategy is not used Zhang and Sutton (2017).
@@ 3.1 Judge Model
Using LLMs to evaluate responses of another LLM, LLM-as-a-judge, is shown promising Zheng et al. (2023); Bai et al. (2022b); Kim et al. (2023). Inspired by these studies, we leverage the generative pairwise evaluator, which we refer to as Judge Model (JM), for aligning LLMs with human preference Ethayarajh et al. (2023); Zhao et al. (2023); Liu et al. (2023). Unlike an RM, which produces a scalar score for a single response, JM simply chooses the better response between the two responses. Specifically, the JM $\pi$ is trained by maximizing the log-likelihood of the judge token $\mathcal{J}\in\{\mathcal{A},\mathcal{B}\}$ corresponding ground-truth preference label for a given judgment template $\mathcal{C}$. Figure 3 illustrates an example of a judgment template.
Figure 3: An example of a judgment template $\mathcal{C}$. The judgment template asks which of the two responses is better for a given prompt and requests to select the judge token $\mathcal{J}\in\{\mathcal{A},\mathcal{B}\}$ corresponding to the better response. Optionally, a principle for the judgment can be added to the judgment template, and the rationale can be included in the target sequence for training.
@@
One advantage of JM compared to the conventional method with RM is that the JM can adapt the judgment template $\mathcal{C}$ according to multiple aspects of human preferences, reflecting diverse principles Sun et al. (2023); Cui et al. (2023). Also, the generative nature of JM makes it easy to utilize rationale, the justification for the judgment in natural language, for more precise judgments, similar to the observations on instruction-tuning of LLMs Mukherjee et al. (2023). For a given principle $p\in\mathcal{P}$ where the $\mathcal{P}$ is a principle set, we devise different judgment templates $\mathcal{C}_{p}$ for each $p$. Also, we can optionally include a rationale $\mathcal{R}$ in the target sequence for the judgment task when it is available. The JM first produces a judge token $\mathcal{J}\in\{\mathcal{A},\mathcal{B}\}$ which corresponds to the ground-truth label of the better response, then rationale $\mathcal{R}$ is followed. Formally, $\pi(\mathcal{J},\mathcal{R}\mid\mathcal{C}_{p}(x,y_{w},y_{l}))$. As a result, principle-aware judgment can be conducted by simply adjusting the judgment template.
@@
Since JM is trained by JSFT, it can judge its own response by playing the roles of the policy and judge: this is the intuition behind the name Self-Judge. As a policy, the JM first samples responses $y_{a},y_{b}$ for a given $x$. Then, the JM, as a judge, chooses a better response between the two responses, $\pi(\mathcal{J}\mid\mathcal{C}(x,y_{a},y_{b}))$. More precisely, it averages the likelihoods of corresponding judge tokens, $\mathcal{J}$, utilizing a position-swapped judgment template to mitigate position bias Chiang et al. (2023). From the relative likelihoods of these judge tokens for each response, a pseudo-preference triplet label $(x,\hat{y}_{w},\hat{y}_{l})$ for the self-training and self-rejection can be deduced. For the principle-aware judgments, we can first check the winning rate across principles and then the mean likelihood of judge tokens across principles for deciding orders when a tie happens.
@@ 3.2 Self-Training by On-Policy Judgment
Ethayarajh et al. (2023) uses JM’s predicted likelihood of judge tokens as the reward for on-policy evaluation in RLHF framework, comparing with a blank baseline response 222huggingface.co/stanfordnlp/SteamSHP-flan-t5-xl. This approach has a significant limitation; the likelihood inferred by a language model cannot be regarded as confidence without calibration Zhou et al. (2023); Zhu et al. (2023). Conversely, to leverage JMs properly, we adopt objectives that perform optimizations on preference orders since these objectives do not require pointwise scalar scores Rafailov et al. (2023); Zhao et al. (2023); Azar et al. (2023). In our framework, we regard the reference policy $\pi_{\text{ref}}$ as a judge to evaluate the preference order between generated samples from current policy $\pi_{\theta}$, as follows:
@@ 3.3 Self-Rejection by Tournament
However, judging all the possible comparison pairs to get the average winning rate requires $\mathcal{O}(N^{2})$ forward passes for (${N\choose 2}$ comparisons. Thus, we adopt the tournament Zhao et al. (2023); Liu et al. (2023) for rejection sampling at inference time. We construct a tournament tree whose leaf nodes are sampled responses, and the non-leaf nodes are chosen by the winner on judgment between the child nodes. Since the tournament tree has less than $N-1$ non-leaf nodes, we can find the best response by $\mathcal{O}(N)$ forward passes, the same as the Best-of-$N$ sampling with a separated RM.
@@ 4.2 Baselines
We choose DPO Rafailov et al. (2023) as an offline learning baseline and use DPO objective for self-training in Self-Judge. We include RSO Liu et al. (2023) with the DPO objective as an off-policy learning baseline, which utilizes the JM as a separate evaluator. We include baselines that utilize RMs, ReST Gulcehre et al. (2023) and RAFT Dong et al. (2023) for off-policy learning approach and RLHF Ziegler et al. (2020) for on-policy learning approach. We choose Llama-2-7B Touvron et al. (2023) as a base model for all experiments for fair comparisons.
@@ 5.2 Analysis of Judge Models
Table 2: Prediction accuracy on test split of HH-Helpful and winning rate on AlpacaEval using JM as a judge or a policy. JSFT improves not only the performance as a policy but also the performance as a judge of JM.
@@
In Footnote 4, we observe that RLHF with JM as an evaluator results in a lower performance compared to conventional RM producing scalar rewards, which implies that the judge token likelihood obtained from a judgment could not be regarded as a pointwise preference score on the Bradley-Terry model Bradley and Terry (1952). Table 4 shows that on-policy learning outperforms offline and off-policy learning approaches using a single JM on self-training. Therefore, we can conclude that on-policy learning with preference orders, such as DPO objective, is the best strategy for JM in order to leverage JM’s superior performance on judgments compared to the RM.
@@
Table 3: Evaluation results of models trained with different evaluator types for RLHF. With JMs, we regard the likelihood of judge token comparing with chosen response $y_{w}$ as a reward [^4] We found that using the blank response fails on the assessment of response since the reward is always close to 1.. JM’s token likelihoods are not appropriate for the pointwise reward function in RLHF.
@@ 5.3 Effects of Principle and Rationale
Figure 4: Winning rate, average response length, and 4-gram repetition on AlpacaEval according to the number of sampling ($N$) for self-rejection on JM and JM-PR after self-training. Even though LLM-as-a-judge tends to favor verbose responses Zheng et al. (2023), JM-PR reliably improves the winning rate as $N$ increases, with smoother increments of response lengths and lower repetitions compared to JM.
@@ 6 Related Work
From preference orders in the static dataset, DPO Rafailov et al. (2023) optimizes LLMs by implicit rewards without a separated RM. IPO Azar et al. (2023) proposes a modified objective using an unbounded preference mapping function to mitigate overfitting on deterministic preferences in the dataset. PCO Xu et al. (2023) utilizes cringe loss for optimization, which considers the token-level likelihood of rejected samples as contrastive training. SPIN Chen et al. (2024) performs iterative training considering the distribution gap between SFT datasets and the model’s responses as preference orders. Self-Rewarding Language Models Yuan et al. (2024) trains LLMs to generate scores for a given response by chain-of-thought reasoning to construct preference datasets by self-generated responses. All these approaches differ from our work in that they do not perform on-policy learning.
@@
The generative pairwise evaluator, which we refer to as JM, has been utilized in previous approaches to alignment learning. ILF Scheurer et al. (2023) selects the response that reflects human-requested feedback through JM. SLiC-HF Zhao et al. (2023) constructs a static preference dataset with responses obtained from the SFT model ordered by JM. RSO Liu et al. (2023) approximates the optimal policy of the RLHF objective with rejection sampling through JM’s judge token likelihood. OAIF Guo et al. (2024) adopts pre-aligned LLMs to perform the pairwise judgment task for on-policy learning, not by fine-tuning to JM. All of these approaches focus on utilizing a separate evaluator and do not address self-improvement.
@@ 7 Discussion
Concurrent Works
Recent works such as Self-Rewarding Language Models Yuan et al. (2024) and OAIF Guo et al. (2024) have intersections with Self-Judge in aspects of self-training and on-policy learning for pairwise preferences. Self-Rewarding Language Models trains the initial policy to act as a judge, but evaluations are performed through chain-of-thought reasoning, which requires a significant computational cost for evaluation. Therefore, off-policy learning with an iterative training scheme was adopted, which can yield inferior results compared to on-policy learning, as shown in Table 4. OAIF similarly utilizes a large language model as a judge for on-policy learning but employs an already aligned large language model as the judge. This necessitates a superior aligned language model and incurs additional computational resources compared to this work, which performs self-improvement using a single model. Table 7 illustrates the difference between these approaches compared to Self-Judge.
@@
Table 7: Comparison between Self-Judge and recent concurrent works on preference alignment learning. Unlike the others, Self-Judge achieves both self-training without separated evaluators and on-policy learning.
Effects of Parameter Size on JSFT
Gao et al. (2023) demonstrates that scaling of parameter size enhances the robustness of RMs in the sense of proxy for the ground truth reward. Similarly, we believe parameter size might significantly affect JMs since they act as a proxy for pairwise human preferences. Table 8 shows that as the parameter size increases, the performance of JM increases without JSFT, but the performance gain from JSFT grows correspondingly as parameter size increases. This implies that the transition effects of imitation learning to judgment task positively correlate with parameter size. Therefore, Self-Judge has the additional advantage of obtaining a strong evaluator from the overparameterization of the policy model, since JSFT directly leads the scaling of the initial policy to the scaling of the evaluator.
Model JSFT Judge (Accuracy) Policy (Win) Llama-2-7B 66.11 67.84 20.26 Llama-2-13B 66.86 72.13 24.00
Table 8: Prediction accuracy on test split of HH-Helpful and winning rate on AlpacaEval with different sizes of base model’s parameter size for JMs. We use base and online splits of HH-Helpful for JSFT.
Policy Model as a Judge for Self-Training
The reference model is regarded as a judge for the policy model on the self-training by default in Self-Judge. This corresponds to the assumption of a static environment during a single iteration of the self-training stage. On the other hand, Figure 5 shows the policy model’s capacity as a judge remained after the self-training stage. This implies that self-training can be conducted even in a dynamic environment by regarding the policy model as a judge whose judgment capability may be improved during the self-training process. However, we find that the policy model as a judge yields slightly lower performances compared to the reference model, as shown in Table 9. Therefore, we can conclude that the static evaluator has benefits in the performance, while the dynamic evaluator still can be utilized with comparable performance.
Policy Evaluator AlpacaEval (Win) MT-Bench (Score) JM 20.26 4.19 $\pi_{\theta}$ 42.38 4.67 $\pi_{\text{ref}}$ 44.88 4.80
Table 9: Evaluation results of models trained on HH-Helpful with different strategies on a judge for the self-training stage. It shows that slight degeneration happens when utilizing the policy model $\pi_{\theta}$ as a judge compared to the reference model $\pi_{\text{ref}}$ as a judge.
@@ Acknowledgements
This work was partly supported by an IITP grant funded by the Korean Government (MSIT) (No. RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University)) and the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. RS-2024-00354218). The authors thank the members of NAVER Cloud and Yonsei University for their constructive comments. We are also grateful to Seungju Han for valuable discussions and helpful feedback on earlier drafts of this paper.
@@ A.5 Prompt Templates
<|system|>
{system message}
<|user|>
{prompt at turn 1}
<|assistant|>
{response at turn 1}[EOS]
<|user|>
{prompt at turn 2}
<|assistant|>
{response at turn 2}[EOS]
$\cdots$
@@
HH-Helpful
Below are a series of dialogues between a human user and an AI assistant. The AI tries to be helpful, harmless, and honest.
UltraFeedback
Below are a series of dialogues between a human user and an AI assistant. The AI tries to be helpful, honest, truthful and follows the instructions from human.
@@
<|system|>
{default system message}
<|user|>
Which is the more helpful, harmless, and honest assistant from a given context?
Context:
User: {prompt at turn 1}
Assistant: {responses at turn 1}
$\cdots$
User: {prompt at turn N}
Assistant A: {responses A at turn N}
Assistant B: {responses B at turn N}
Please choose either A or B.
<|assistant|>
Sure! The option which is more helpful, harmless, and honest would be Assistant
{A or B}
@@
<|system|>
{default system message}
<|user|>
Which is the better response to be an assistant who is helpful, honest, truthful and following the given instruction from user?
Instruction:
{prompt}
Response A:
{response A}
Response B:
{response B}
Please choose either A or B.
<|assistant|>
Sure! The option which is more helpful, honest, truthful and following the given instruction from user would be Response
{A or B}
@@
Helpfulness
Under the principle of ’helpfulness’, the assistant should provide users with accurate, relevant, and up-to-date information, ensuring that the content is positive, interesting, engaging, educational, and helpful.
Honesty
Under the principle of ’honesty’, the assistant should be honest about whether it knows the answer and express its uncertainty explicitly. The assistant should confident on questions it knows well and be modest on those it is unfamiliar with using weakeners such as ’I guess’, ’I suppose’, ’probably’, and ’perhaps’ to express uncertainty.
Instruction Following
Under the principle of ’instruction following’, the assistant should align the output with intent of instruction, by understanding the task goal (intended outcome) and restrictions (text styles, format or designated methods, etc.).
Truthfulness
Under the principle of ’truthfulness’, the assistant should answer truthfully and be faithful to factual knowledge as well as given contexts, never making up any new facts that aren’t true or cannot be grounded in the instruction.
@@
<|system|>
{principle-aware system message}
<|user|>
Which is the better response for an assistant when following the principle of ’{principle}’ for a given instruction?
Instruction:
{prompt}
Response A:
{response A}
Response B:
{response B}
Please choose either A or B according to the principle of ’{principle}’.
<|assistant|>
Sure! The option which is better guided by the principle of ’{principle}’ would be Response
{A or B}
@@
<|system|>
{principle-aware system message}
<|user|>
Which is the better response for an assistant when following the principle of ’{principle}’ for a given instruction?
Instruction:
{prompt}
Response A:
{response A}
Response B:
{response B}
Please choose either A or B and evaluate each response according to the principle of ’{principle}’.
<|assistant|>
Sure! Here is the decision of better response and evaluations of each response, guided by the principle of ’{principle}’.
Decision: Response
{A or B}
Evaluation of Response A: {rationale A}
Evaluation of Response B: {rationale B}